OpenStudy (anonymous):

Which of the following sums would be under the radical symbol to find the distance between the points (7, -1) and (-8, -9)? 1^2 + 10^2 1^2 + 8^2 15^2 + 8^2

zepdrix (zepdrix):

\[\Large\bf\sf (\color{#8C3D62}{x_1},\;\color{#F35633}{y_1})=(\color{#8C3D62}{7},\;\color{#F35633}{-1}),\qquad\qquad (\color{royalblue}{x_2},\;\color{#CC0033}{y_2})=(\color{royalblue}{-8},\;\color{#CC0033}{-9})\] Here is our distance formula: \[\Large\bf\sf d\quad=\quad \sqrt{(\color{royalblue}{x_2}-\color{#8C3D62}{x_1})^2+(\color{#CC0033}{y_2}-\color{#F35633}{y_1})^2}\]Understand how to plug the values in? Match up the colors! :)

OpenStudy (anonymous):

It's x1-x2 and y1-y2 though

zepdrix (zepdrix):

We're squaring the quantities, so it won't matter. (x2-x1) is simply the negative of (x1-x2). And when we square it, the negative goes away. But ya if you remember the formula with x1 first, then my bad ^^ might look a little confusing.

OpenStudy (anonymous):

Lol yeah.. I've tried setting it up, but I didn't understand the answer

zepdrix (zepdrix):

\[\Large\bf\sf d\quad=\quad \sqrt{(\color{royalblue}{x_2}-\color{#8C3D62}{x_1})^2+(\color{#CC0033}{y_2}-\color{#F35633}{y_1})^2}\]Plug in the numbers!! +_+\[\Large\bf\sf d\quad=\quad \sqrt{(\color{royalblue}{-8}-\color{#8C3D62}{7})^2+(\color{#CC0033}{-9}-(\color{#F35633}{-1}))^2}\]

zepdrix (zepdrix):

understand how to deal with the -(-1)?

OpenStudy (anonymous):

Turn it into a positive 1?

OpenStudy (anonymous):

Is the answer 1^2+8^2 ?

zepdrix (zepdrix):

Your second value looks good, 8^2. You added the 1 to -9, gives you (-8)^2 which is 8^2. Good good.

zepdrix (zepdrix):

The first value is -8-7. That does not give us 1 or negative 1. Hmmmm check your subtraction again! :O

OpenStudy (anonymous):

Can you help me with this one too? Find the distance between the points (2.7, 5.1) and (3, 4.7). 0.05 0.5 5

zepdrix (zepdrix):

\[\Large\bf\sf d\quad=\quad \sqrt{(\color{royalblue}{3}-\color{#8C3D62}{2.7})^2+(\color{#CC0033}{4.7}-\color{#F35633}{5.1})^2}\]

zepdrix (zepdrix):

Simplifyyyy \c:/ What do you get for your squares?